### v1

# MagicDrive: Street View Generation with Diverse 3D Geometry Control

Abstract
Recent advancements in diffusion models have significantly enhanced the data synthesis with 2D control. Yet, precise 3D control in street view generation, crucial for 3D perception tasks, remains elusive. Specifically, utilizing Bird’s-Eye View (BEV) as the primary condition often leads to challenges in geometry control (e.g., height), affecting the representation of object shapes, occlusion patterns, and road surface elevations, all of which are essential to perception data synthesis, especially for 3D object detection tasks. In this paper, we introduce MagicDrive, a novel street view generation framework offering diverse 3D geometry controls, including camera poses, road maps, and 3D bounding boxes, together with textual descriptions, achieved through tailored encoding strategies. Besides, our design incorporates a cross-view attention module, ensuring consistency across multiple camera views. With MagicDrive, we achieve high-fidelity street-view synthesis that captures nuanced 3D geometry and various scene descriptions, enhancing tasks like BEV segmentation and 3D object detection. 00footnotetext: $*{}^{*}$qual contribution. ${}^{{\dagger}}$orresponding authors. Project Page: https://flymin.github.io/magicdrive.
Figure 1: Multi-camera street view generation from MagicDrive. MagicDrive can generate continuous camera views with controls from road map, object boxes, and text (e.g., weather).
1 Introduction
The high costs associated with data collection and annotation often impede the effective training of deep learning models. Fortunately, cutting-edge generative models have illustrated that synthetic data can notably boost performance across various tasks, such as object detection (Chen et al., 2023b) and semantic segmentation (Wu et al., 2023). Yet, the prevailing methodologies are largely tailored to 2D contexts, primarily relying on 2D bounding boxes (Lin et al., 2014; Han et al., 2021) or segmentation maps (Zhou et al., 2019) as layout conditions (Chen et al., 2023b; Li et al., 2023).
In autonomous driving applications, a thorough grasp of the 3D environment is essential. This demands reliable techniques for tasks like Bird’s-Eye View (BEV) map segmentation (Zhou & Krähenbühl, 2022; Ji et al., 2023) and 3D object detection (Chen et al., 2020; Huang et al., 2021; Liu et al., 2023; Ge et al., 2023). A genuine 3D geometry representation is crucial for capturing intricate details from 3D annotations, such as road elevations, object heights, and their occlusion patterns, as shown in Figure 2. Consequently, generating multi-camera street-view images according to 3D annotations becomes vital to boost downstream perception tasks.
Figure 2: 3D bounding boxes are crucial for street view synthesis. Two examples show that 2D boxes or BEV maps lost distance, height, and elevation. Images are generated from MagicDrive.
For street-view data synthesis, two pivotal criteria are realism and controllability. Realism requires that the quality of the synthetic data should align with that of real data; and in a given scene, views from varying camera perspectives should remain consistent with one another (Mildenhall et al., 2020). On the other hand, controllability emphasizes the precision in generating street-view images that adhere to provided conditions: the BEV map, 3D object bounding boxes, and camera poses for views. Beyond these core requirements, effective data augmentation should also grant the flexibility to tweak finer scenario attributes, such as prevailing weather conditions or the time of day. Existing solutions like BEVGen (Swerdlow et al., 2023) approach street view generation by encapsulating all semantics within BEV. Conversely, BEVControl (Yang et al., 2023) starts by projecting 3D coordinates to image views, subsequently using 2D geometric guidance. However, both methods compromise certain geometric dimensions—height is lost in BEVGen and depth in BEVControl.
The rise of diffusion models has significantly pushed the boundaries of controllable image generation quality. Specifically, ControlNet (Zhang et al., 2023a) proposes a flexible framework to incorporate 2D spatial controls based on pre-trained Text-to-Image (T2I) diffusion models (Rombach et al., 2022). However, the challenge of seamlessly integrating multiple 3D conditions with multi-camera view consistency in street view synthesis remains unresolved.
In this paper, we introduce MagicDrive, a novel framework dedicated to street-view synthesis with diverse 3D geometry controls. For realism, we harness the power of pre-trained stable diffusion (Rombach et al., 2022), further fine-tuning it for street view generation. One distinctive component of our framework is the cross-view attention module. This simple yet effective component provides multi-view consistency through interactions between adjacent views. In contrast to previous methods, MagicDrive encodes objects and road maps separately, amplifying the utilization of all available 3D data for controllability enhancement. More specifically, given the sequence-like, variable-length nature of 3D bounding boxes, we employ cross-attention akin to text embeddings for their encoding. Besides, we adapt ControlNet (Zhang et al., 2023a) to view transformation and incorporate road map information. Without resorting to explicit geometric transformations or imposing geometric constraints on multi-camera consistency, our design learns the inherent geometric relationships from data, significantly simplifying our framework. Finally, MagicDrive factors in textual descriptions, offering attribute control such as weather conditions and time of day.
Our MagicDrive framework, despite its simplicity, excels in generating strikingly realistic images that align with road maps, 3D bounding boxes, and varied camera perspectives. Besides, the images produced can enhance the training for both 3D object detection and BEV segmentation tasks. Furthermore, MagicDrive offers comprehensive geometric controls at the scene, background, and foreground levels. This flexibility makes it possible to craft previously unseen street views suitable for simulation purposes. We summarize the main contributions of this work as:
  • The introduction of MagicDrive, an innovative framework that generates multi-perspective camera views conditioned on BEV and 3D data tailored for autonomous driving.
  • The development of simple yet potent strategies to manage 3D geometric data, effectively addressing the challenges of multi-camera view consistency.
  • Through rigorous experiments, we demonstrate that MagicDrive outperforms prior street view generation techniques, notably for the multi-dimensional controllability. Additionally, our results reveal that synthetic data delivers considerable improvements in 3D perception tasks.
2 Related Work
Diffusion Models for Conditional Generation. Diffusion models (Ho et al., 2020; Song et al., 2020) generate images by learning a progressive denoising process from the Gaussian noise distribution to the image distribution. These models have proven exceptional across diverse tasks, such as text-to-image synthesis (Rombach et al., 2022; Nichol et al., 2022), inpainting (Wang et al., 2023), and instructional image editing (Zhang et al., 2023b; Brooks et al., 2023), due to their adaptability and competence in managing various form of controls (Rombach et al., 2022; Zhang et al., 2023a; Li et al., 2023) and multiple conditions (Liu et al., 2022a; Huang et al., 2023; Gao et al., 2023). Besides, data synthesized from geometric annotations can aid downstream tasks such as 2D object detection (Chen et al., 2023b; Wu et al., 2023). Thus, this paper explores the potential of T2I diffusion models in generating street-view images and benefiting downstream 3D perception models.
Street View Generation. Numerous street view generation models condition on 2D layouts, such as 2D bounding boxes (Li et al., 2023) and semantic segmentation (Wang et al., 2022). These methods leverage 2D layout information corresponding directly to image scale, whereas the 3D information does not possess this property, thereby rendering such methods unsuitable for leveraging 3D information for generation. For street view synthesis with 3D geometry, BEVGen (Swerdlow et al., 2023) is the first to explore. It utilizes a BEV map as a condition for both roads and vehicles. However, the omission of height information limits its application in 3D object detection. BEVControl (Yang et al., 2023) amends the loss of object’s height by the height lifting process, but the projection from 3D to 2D results in the loss of essential 3D geometric information, like depth and occlusion. Thus, neither of them fully exploits 3D annotations or can utilize textual control of driving scenes. In this paper, we propose to encode bounding boxes and road maps separately for more nuanced control and integrate scene descriptions, offering enhanced control over the generation of street views.
Multi-camera Image Generation of a 3D scene fundamentally requires viewpoint consistency. Several studies have addressed this issue within the context of indoor scenes. For instance, MVDiffusion (Tang et al., 2023) employs panoramic images and a cross-view attention module to maintain global consistency, while Tseng et al. (2023) leverage epipolar geometry as a constraining prior. These approaches, however, primarily rely on the continuity of image views, a condition not always met in street views due to limited camera overlap. Our MagicDrive introduces extra cross-view attention modules to UNet, which significantly enhances consistency across multi-camera views.
3 Preliminary
Problem Formulation. In this paper, we consider the coordinate of the LiDAR system as the ego car’s coordinate, and parameterize all geometric information according to it. Let $\mathbf{S}=\{\mathbf{M},\mathbf{B},\mathbf{L}\}$ be the description of a driving scene around the ego vehicle, where $\mathbf{M}\in\{0,1\}^{w\times h\times c}$ is the binary map representing a $w\times h$ meter road area in BEV with $c$ semantic classes, $\mathbf{B}=\{(c_{i},b_{i})\}_{i=1}^{N}$ represents the 3D bounding box position ($b_{i}=\{(x_{j},y_{j},z_{j})\}_{j=1}^{8}\in\mathbb{R}^{8\times 3}$) and class ($c_{i}\in\mathcal{C}$) for each object in the scene, and $\mathbf{L}$ is the text describing additional information about the scene (e.g., weather and time of day). Given a camera pose $\mathbf{P}=[\mathbf{K},\mathbf{R},\mathbf{T}]$ (i.e., intrinsics, rotation, and translation), the goal of street-view image generation is to learn a generator $\mathcal{G}(\cdot)$ which synthesizes realistic images $I\in\mathbb{R}^{H\times W\times 3}$ corresponding to the scene $\mathbf{S}$ and camera pose $\mathbf{P}$ as, $I=\mathcal{G}(\mathbf{S},\mathbf{P},z)$, where $z\sim\mathcal{N}(0,1)$ is a random noise from Gaussian distribution.
Conditional Diffusion Models. Diffusion models (Ho et al., 2020; Song et al., 2020) generate data ($\bm{x}_{0}$) by iteratively denoising a random Gaussian noise ($\bm{x}_{T}$) for $T$ steps. Typically, to learn the denoising process, the network is trained to predict the noise by minimizing the mean-square error:
$\ell_{simple}=\mathbb{E}_{\bm{x}_{0},\bm{c},\bm{\epsilon},t}\left[||\bm{\epsilon}-\bm{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},t,\bm{c})||^{2}\right]\text{,}$ (1)
where $\bm{\epsilon}_{\theta}$ is the network to train, with parameters $\theta$, $\bm{c}$ is optional conditions, which is used for the conditional generation, $t\in[0,T]$ is the time-step, $\bm{\epsilon}\in\mathcal{N}(0,I)$ is the additive Gaussian noise, and $\bar{\alpha}_{t}$ is a scalar parameter. Latent diffusion models (LDM) (Rombach et al., 2022) is a special kind of diffusion model, where they utilize a pre-trained Vector Quantized Variational AutoEncoder (VQ-VAE) (Esser et al., 2021) and perform diffusion process in the latent space. Given the VQ-VAE encoder as $z=\mathcal{E}(x)$, one can rewrite $\bm{\epsilon}_{\theta}(\cdot)$ in Equation 1 as $\bm{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{t}}\mathcal{E}(\bm{x}_{0})+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},t,\bm{c})$ for LDM. Besides, LDM considers text describing the image as condition $c$.
Figure 3: Overview of MagicDrive for street-view image generation. MagicDrive generates highly realistic images, exploiting geometric information from 3D annotations by independently encoding road maps, object boxes, and camera parameters for precise, geometry-guided synthesis. Additionally, MagicDrive accommodates guidance from descriptive conditions (e.g., weather).
4 Street View Generation with 3D Information
The overview of MagicDrive is depicted in Figure 3. Operating on the LDM pipeline, MagicDrive generates street-view images conditioned on both scene annotations ($\mathbf{S}$) and the camera pose ($\mathbf{P}$) for each view. Given the 3D geometric information in scene annotations, projecting all to a BEV map, akin to BEVGen (Swerdlow et al., 2023) or BEVControl (Yang et al., 2023), doesn’t ensure precise guidance for street view generation, as exemplified in Figure 2. Consequently, MagicDrive categorizes conditions into three levels: scene (text and camera pose), foreground (3D bounding boxes), and background (road map); and integrates them separately via cross-attention and an additive encoder branch, detailed in Section 4.1. Additionally, maintaining consistency across different cameras is crucial for synthesizing street views. Thus, we introduce a simple yet effective cross-view attention module in Section 4.2. Lastly, we elucidate our training strategies in Section 4.3, emphasizing Classifier-Free Guidance (CFG) in integrating various conditions.
4.1 Geometric Conditions Encoding
As illustrated in Figure 3, two strategies are employed for information injection into the UNet of diffusion models: cross-attention and additive encoder branch. Given that the attention mechanism (Vaswani et al., 2017) is tailored for sequential data, cross-attention is apt for managing variable length inputs like text tokens and bounding boxes. Conversely, for grid-like data, such as road maps, the additive encoder branch is effective in information injection (Zhang et al., 2023a). Therefore, MagicDrive employs distinct encoding modules for various conditions.
Scene-level Encoding includes camera pose $\mathbf{P}=\{\mathbf{K}\in\mathbb{R}^{3\times 3},\mathbf{R}\in\mathbb{R}^{3\times 3},\mathbf{T}\in\mathbb{R}^{3\times 1}\}$, and text sequence $\mathbf{L}$. For text, we construct the prompt with a template as “A driving scene image at {location}. {description} ”, and leverage a pre-trained CLIP text encoder ($E_{text}$) as LDM (Rombach et al., 2022), as shown by Equation 2, where $L$ is the token length of $\mathbf{L}$. As for camera pose, we first concat each parameter by their column, resulting in $\bar{\mathbf{P}}=[\mathbf{K},\mathbf{R},\mathbf{T}]^{T}\in\mathbb{R}^{7\times 3}$. Since $\bar{\mathbf{P}}$ contains values from $\sin$/$\cos$ functions and also 3D offsets, to have the model effectively interpret these high-frequency variations, we apply Fourier embedding (Mildenhall et al., 2020) to each 3-dim vector before leveraging a Multi-Layer Perception (MLP, $E_{cam}$) to embed the camera pose parameters, as in Equation 3. To maintain consistency, we set the dimension of $h^{c}$ the same as that of ${h^{t}_{i}}$. Through the CLIP text encoder, each text embedding $h^{t}_{i}$ already contains positional information (Radford et al., 2021). Therefore, we prepend the camera pose embedding $h^{c}$ to text embeddings, resulting in scene-level embedding $\bm{h}^{s}=[h^{c},\bm{h}^{t}]$.
$\bm{h}^{t}$ $=[h^{t}_{1}\dots h^{t}_{L}]=E_{text}(\mathbf{L}),$ (2)
$h^{c}$ $=E_{cam}(\operatorname{Fourier}(\bar{\mathbf{P}}))=E_{cam}(\operatorname{Fourier}([\mathbf{K},\mathbf{R},\mathbf{T}]^{T})).$ (3)
3D Bounding Box Encoding. Since each driving scene has a variable length of bounding boxes, we inject them through the cross-attention mechanism similar to scene-level information. Specifically, we encode each box into a hidden vector $h^{b}$, which has the same dimensions as that of $h^{t}$. Each 3D bounding box $(c_{i},b_{i})$ contains two types of information: class label $c_{i}$ and box position $b_{i}$. For class labels, we utilize the method similar to Li et al. (2023), where the pooled embeddings of class names ($L_{c_{i}}$) are considered as label embeddings. For box positions $b_{i}\in\mathbb{R}^{8\times 3}$, represented by the coordinates of its 8 corner points, we utilize Fourier embedding to each point and pass through an MLP for encoding, as in Equation 4. Then, we use an MLP to compress both class and position embedding into one hidden vector, as in Equation 5. The final hidden states for all bounding boxes of each scene are represented as $\bm{h}^{b}=[h^{b}_{1}\dots h^{b}_{N}]$, where $N$ is the number of boxes.
$e^{b}_{c}(i)=\operatorname{AvgPool}(E_{text}(L_{c_{i}})),\;e^{b}_{p}(i)=\operatorname{MLP}_{p}(\operatorname{Fourier}(b_{i})),$ (4)
$h^{b}_{i}=E_{box}(c_{i},b_{i})=\operatorname{MLP}_{b}(e^{b}_{c}(i),e^{b}_{p}(i)).$ (5)
Ideally, the model learns the geometric relationship between bounding boxes and camera pose through training. However, the distribution of the number of visible boxes to different views is long-tailed. Thus, we bootstrap learning by filtering visible objects to each view ($v_{i}$), i.e., $f_{viz}$ in Equation 6. Besides, we also add invisible boxes for augmentation (more details in Section 4.3).
$\bm{h}^{b}_{v_{i}}=\{h^{b}_{i}\in\bm{h}^{b}|f_{viz}(b_{i},\mathbf{R}_{v_{i}},\mathbf{T}_{v_{i}})>0\}.$ (6)
Road Map Encoding. The road map has a 2D-grid format. While Zhang et al. (2023a) shows the addictive encoder can incorporate this kind of data for 2D guidance, the inherent perspective differences between the road map’s BEV and the camera’s First-Person View (FPV) create discrepancies. BEVControl (Yang et al., 2023) employs a back-projection to transform from BEV to FPV but complicates the situation with an ill-posed problem. In MagicDrive, we propose that explicit view transformation is unnecessary, as sufficient 3D cues (e.g., height from object boxes and camera pose) allow the addictive encoder to accomplish view transformation. Specifically, we integrate scene-level and 3D bounding box embeddings into the map encoder (see Figure 3). Scene-level embeddings provide camera poses, and box embeddings offer road elevation cues. Additionally, incorporating text descriptions facilitates the generation of roads under varying conditions (e.g., weather and time of day). Thus, the map encoder can synergize with other conditions for generation.
Figure 4: Cross-view Attention. left: we introduce cross-view attention to the pre-trained UNet after the cross-attention module. right: we highlight some areas for comparison between without and with cross-view attention. Cross-view attention guarantees consistency across multiple views.
4.2 Cross-view Attention Module
In multi-camera view generation, it is crucial that image synthesis remains consistent across different perspectives. To maintain consistency, we introduce a cross-view attention module (Figure 4). Given the sparse arrangement of cameras in driving contexts, each cross-view attention allows the target view to access information from its immediate left and right views, as in Equation 8; here, $t$, $l$, and $r$ are the target, left, and right view respectively. Then, the target view aggregates such information with skip connection, as in Equation 9, where $\bm{h}^{v}$ indicates the hidden state of the target view.
$\operatorname{Attention}_{cv}^{i}(Q_{t},K_{i},V_{i})=\operatorname{softmax}(\begin{matrix}\frac{Q_{t}K_{i}^{T}}{\sqrt{d}}\end{matrix})\cdot V_{i},\;i\in\{l,r\}\text{,}$ (8)
$\bm{h}_{out}^{v}=\bm{h}_{in}^{v}+\operatorname{Attention}_{cv}^{l}+\operatorname{Attention}_{cv}^{r}\text{.}$ (9)
We inject cross-view attention after the cross-attention module in the UNet and apply zero-initialization (Zhang et al., 2023a) to bootstrap the optimization. The efficacy of the cross-view attention module is demonstrated in Figure 4 right, Figure 5, and Figure 6. The multilayered structure of UNet enables aggregating information from long-range views after several stacked blocks. Therefore, using cross-view attention on adjacent views is enough for multi-view consistency, further evidenced by the ablation study in Appendix C.
4.3 Model Training
Classifier-free Guidance reinforces the impact of conditional guidance (Ho & Salimans, 2021; Rombach et al., 2022). For effective CFG, models need to discard conditions during training occasionally. Given the unique nature of each condition, applying a drop strategy is complex for multiple conditions. Therefore, our MagicDrive simplifies this for four conditions by concurrently dropping scene-level conditions (camera pose and text embeddings) at a rate of $\gamma^{s}$. For boxes and maps, which have semantic representations for null (i.e., padding token in boxes and 00 in maps) in their encoding, we maintain them throughout training. At inference, we utilize null for all conditions, enabling meaningful amplification to guide generation.
Training Objective and Augmentation. With all the conditions injected as inputs, we adapt the training objective described in Section 3 to the multi-condition scenario, as in Equation 10.
$\ell=\mathbb{E}_{\bm{x}_{0},\bm{\epsilon},t,\{\bm{S},\bm{P}\}}\left[||\bm{\epsilon}-\bm{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{t}}\mathcal{E}(\bm{x}_{0})+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},t,\{\mathbf{S},\mathbf{P}\})\right].$ (10)
Besides, we emphasize two essential strategies when training our MagicDrive. First, to counteract our filtering of visible boxes, we randomly add 10% invisible boxes as an augmentation, enhancing the model’s geometric transformation capabilities. Second, to leverage cross-view attention, which facilitates information sharing across multiple views, we apply unique noises to different views in each training step, preventing trivial solutions to Equation 10 (e.g., outputting the shared component across different views). Identical random noise is reserved exclusively for inference.
5 Experiments
5.1 Experimental Setups
Dataset and Baselines. We employ the nuScenes dataset (Caesar et al., 2020) as the testing ground for MagicDrive, a prevalent dataset in BEV segmentation and detection for driving. We adhere to the official configuration, utilizing 700 street-view scenes for training and 150 for validation. Our baselines are BEVGen (Swerdlow et al., 2023) and BEVControl (Yang et al., 2023), both recent propositions for street view generation. Our method considers 10 object classes and 8 road classes, surpassing the baseline models in diversity. Appendix B holds additional details.
Evaluation Metrics. We evaluate both realism and controllability for street view generation. Realism is mainly measured using Fréchet Inception Distance (FID), reflecting image synthesis quality. For controllability, MagicDrive is evaluated through two perception tasks: BEV segmentation and 3D object detection, with CVT (Zhou & Krähenbühl, 2022) and BEVFusion (Liu et al., 2023) as perception models, respectively. Both of them are renowned for their performance in each task. Firstly, we generate images aligned with the validation set annotations and use perception models pre-trained with real data to assess image quality and control accuracy. Then, data is generated based on the training set to examine the support for training perception models as data augmentation.
Model Setup. Our MagicDrive utilizes pre-trained weights from Stable Diffusion v1.5, training only newly added parameters. Per Zhang et al. (2023a), a trainable UNet encoder is created for $E_{map}$. New parameters, except for the zero-init module and the class token, are randomly initialized. We adopt two resolutions to reconcile discrepancies in perception tasks and baselines: 224$\times$400 (0.25$\times$ down-sample) following BEVGen and for CVT model support, and a higher 272$\times$736 (0.5$\times$ down-sample) for BEVFusion support. Unless stated otherwise, images are sampled using the UniPC (Zhao et al., 2023) scheduler for 20 steps with CFG at $2.0$.
Table 1: Comparison of generation fidelity with driving-view generation methods. Conditions for data synthesis are from nuScenes validation set. For each task, we test the corresponding models trained on the nuScenes training set. MagicDrive surpasses all baselines throughout the evaluation. $\uparrow$/$\downarrow$ indicates that a higher/lower value is better. The best results are in bold, while the second best results are in underlined italic (when other methods are available).
[…]
Road mIoU $\uparrow$ Vehicle mIoU $\uparrow$ mAP $\uparrow$ NDS $\uparrow$
Oracle - - 72.21 33.66 35.54 41.21
BEVGen 224$\times$400 25.54 50.20 5.89 - -
BEVControl - 24.85 60.80 26.80 - -
MagicDrive 224$\times$400 16.20 61.05 27.01 12.30 23.32
MagicDrive 272$\times$736 16.59 54.24 31.05 20.85 30.26
Figure 5: Qualitative comparison with baselines. Both scenes are from nuScenes validation set. We highlight some areas with rectangles to ease comparison. Up: compared with BEVControl, generations from MagicDrive appear more consistent in both background and foreground. Down: compared with BEVControl, image quality of objects from MagicDrive is much better.
Table 2: Comparison about support for 3D object detection model (i.e., BEVFusion). MagicDrive generates 272×736 images for augmentation. Results are reported on the nuScenes validation set. Modality Data mAP $\uparrow$ NDS $\uparrow$ C w/o synthetic data 32.48 37.61 w/ MagicDrive 35.14 +2.66 39.63 +2.02 C+L w/o synthetic data 64.92 69.42 w/ MagicDrive 67.28 +2.36 70.14 +0.72 Table 3: Comparison about support for BEV segmentation model (i.e., CVT). Results are reported by testing on the nuScenes validation set. Data Vehicle mIoU $\uparrow$ Road mIoU $\uparrow$ w/o synthetic data 36.00 74.30 w/ BEVGen 36.60 +0.60 71.90 -2.40 w/ MagicDrive 40.34 +4.34 79.56 +5.26
[Uncaptioned image] Figure 6: Showcase for multi-level control with MagicDrive. We show scene-level, background-level, and foreground-level control separately with diffusion conditions. All scenes are based on the nuScenes validation set.
5.2 Main Results
Realism and Controllability Validation. To confirm the efficacy of MagicDrive in generating realistic images, we employ the nuScenes validation set to synthesize street-view images and show the metrics in Table 1. MagicDrive surpasses all other methods in image quality, delivering significantly lower FID scores. Despite the inherent domain gap between stable diffusion pre-training and street views, our refinement process effectively mitigates this, yielding highly convincing results. Regarding controllability, assessed via BEV segmentation tasks, MagicDrive consistently equals or exceeds baseline results at 224$\times$400 resolution. This is mainly attributed to the distinct encoding design which significantly boosts generation precision on vehicles. At 272$\times$736 resolution, advancements in our encoding strategy are demonstrated by enhanced vehicle mIoU performance. Cropping large areas of each view affects the road mIoU on CVT negatively. Nonetheless, BEVFusion’s results in 3D object detection further corroborate the efficacy of our bounding box encoding.
Training Support for BEV Segmentation and 3D Object Detection. MagicDrive can produce augmented data with accurate annotation controls, enhancing the training for perception tasks. For BEV segmentation, we augment an equal number of images as in the original dataset, ensuring consistent training iterations and batch sizes for fair comparisons to the baseline. As shown in Table 3, MagicDrive significantly enhances CVT in both settings, outperforming BEVGen, which only marginally improves vehicle segmentation. For 3D object detection, we train BEVFusion models with MagicDrive ’s synthetic data as augmentation. To optimize data augmentation, we randomly exclude 50% of bounding boxes in each generated scene. Table 2 shows the advantageous impact of MagicDrive ’s data in both CAM-only (C) and CAM+LiDAR (C+L) settings. It’s crucial to note that in CAM+LiDAR settings, BEVFusion utilizes both modalities for object detection, requiring more precise image generation due to LiDAR data incorporation. Nevertheless, MagicDrive’s synthetic data integrates seamlessly with LiDAR inputs, highlighting the data’s high fidelity.
5.3 Qualitative Evaluation
Comparison with Baselines. We assessed MagicDrive against two baselines, BEVGen and BEVControl, synthesizing multi-camera views for the same validation scenes. Figure 5 illustrates that MagicDrive generates images markedly superior in quality to both baselines, particularly excelling in accurate object positioning and maintaining consistency in street views for backgrounds and objects, notably over BEVControl. Such performance primarily stems from MagicDrive ’s bounding box encoder and its cross-view attention module.
Multi-level Controls. The design of MagicDrive introduces multi-level controls to street-view generation through separation encoding. This section demonstrates the capabilities of MagicDrive by exploring three control signal levels: scene level (time of day and weather), background level (BEV map alterations and conditional views), and foreground level (object orientation and deletion). As illustrated in Figure 1 and Figure 6, MagicDrive adeptly accommodates alterations at each control level, maintaining multi-camera consistency and high realism in generation.
6 Ablation Study
[…]
Bounding Box Encoding. MagicDrive utilizes separate encoders for bounding boxes and road maps. To demonstrate the efficacy, we train a ControlNet (Zhang et al., 2023a) that takes the BEV map with both road and object semantics as a condition (like BEVGen), denoted as “w/o $E_{box}$” in Table 4. Given the relatively small size of objects in the BEV map, it is necessary for a separate $E_{box}$ to effectively help the model accurately represent vehicle annotations, evidenced by the performance gap in vehicle mIoU. Besides, by applying visible object filter $f_{viz}$, both the road and the vehicle mIoU improve significantly. This is due to that visibility filtering reduces the burden of optimization. Additionally, we explored a variant of MagicDrive, incorporating $E_{box}$ alongside BEV containing both road and object semantics. However, this redundancy did not enhance performance, reinforcing the merit of integrating diverse types of information through different strategies.
Effect of Classifier-free Guidance. We focus on the two most crucial conditions, i.e. object boxes and road maps, and analyze how CFG affects the performance of generation. We change CFG from 1.5 to 4.0 and plot the change of validation results from CVT in Figure 7. Firstly, by increasing the CFG scale, FID keeps degrading, since CFG significantly changes the contrast and sharpness, which is also observed by previous works (Chen et al., 2023b). Secondly, by keeping the map the same for both conditional inference and unconditional inference, we can eliminate the effect of CFG on the map condition. As shown by blue lines of Figure 7, increasing CFG scale results in the highest vehicle mIoU at CFG=2.5, but the road mIoU keeps decreasing. Thirdly, with $M=\{0\}$ for unconditional inference in CFG, road mIoU significantly increases. However, it slightly degrades the guidance on vehicle generation. As stated in Section 4.3, CFG gains complication through incorporating more conditions. Although we ease the case for training, we still face different choices of CFG during inferece. We leave the in-depth investigation for this case as future work.
7 Conclusion
This paper presents MagicDrive, a novel framework to encode multiple geometric controls for high-quality multi-camera street view generation. With the separation encoding design, MagicDrive fully utilizes geometric information from 3D annotations and realizes accurate semantic control for street views. Besides, the proposed cross-view attention module is simple yet effective in guaranteeing consistency across multi-camera views. As evidenced by experiments, the generations from MagicDrive show high realism and fidelity to 3D annotations. Multiple controls equipped MagicDrive with improved generalizability for the generation of novel street views. Meanwhile, MagicDrive can be used for data augmentation, facilitating the training for perception models on both BEV segmentation and 3D object detection tasks.
Figure 8: Failure cases of MagicDrive.
Limitation and Future Work. We show failure cases from MagicDrive in Figure 8. Although MagicDrive can generate night views, they are not as dark as real images (as in Figure 8a). This may be due to that diffusion models are hard to generate too dark images (Guttenberg, 2023). Figure 8b shows that MagicDrive cannot generate unseen weathers for nuScenes. Future work may focus on how to improve the cross-domain generalization ability of street view generation.
Acknowledgement. This work is supported in part by General Research Fund (GRF) of Hong Kong Research Grants Council (RGC) under Grant No. 14203521 and in part by the Research Matching Grant Scheme under Grant No. 8601109. We gratefully acknowledge the support of MindSpore, CANN (Compute Architecture for Neural Networks) and Ascend AI Processor used for this research.
APPENDIX
Appendix A Object Filtering
Figure 9: Bounding box filtering to each view. The dashed lines in (b-c) represent the $x$-axis of the camera’s coordinates. Boxes are retained only if they have at least a point in the positive half of the $z$-axis in each camera’s coordinate.
In Equation 6, we employ $f_{viz}$ for object filtering to facilitate bootstrap learning. We show more details of $f_{viz}$ here. Refer to Figure 9 for illustration. For the sake of simplicity, each camera’s Field Of View (FOV) is not considered. Objects are defined as visible if any corner of their bounding boxes is located in front of the camera (i.e., $z^{v_{i}}>0$) within each camera’s coordinate system. The application of $f_{viz}$ significantly lightens the workload of the bounding box encoder, evidence for which can be found in Section 6.
Appendix B More Experimental Details
Semantic Classes for Generation. To support most perception models on nuScenes, we try to include semantics commonly used in most settings (Huang et al., 2021; Zhou & Krähenbühl, 2022; Liu et al., 2023; Ge et al., 2023). Specifically, for objects, ten categories include car, bus, truck, trailer, motorcycle, bicycle, construction vehicle, pedestrian, barrier, and traffic cone. For the road map, eight categories include drivable area, pedestrian crossing, walkway, stop line, car parking area, road divider, lane divider, and roadblock.
Optimization. We train all newly added parameters using AdamW (Loshchilov & Hutter, 2019) optimizer and a constant learning rate at $8e^{-5}$ and batch size 24 (total 144 images for 6 views) with a linear warm-up of 3000 iterations, and set $\gamma^{s}=0.2$.
Appendix C Ablation on Number of Attending Views
Table 5: Ablation on number of attending views. Evaluation results are from CVT on the synthetic nuScenes validation set, without $M=\{0\}$ in CFG scale $=2$.
Attending Num. FID $\downarrow$ Road mIoU $\uparrow$ Vehicle mIoU $\uparrow$
1 13.06 58.63 26.41
2 (ours) 14.46 59.31 27.13
5 (all) 14.76 58.35 25.41
Figure 10: Comparison between different numbers of attending views. Only attending to one view results in worse multi-camera consistency.
In Table 5, we demonstrate the impact of varying the number of attended views on evaluation results. Attending to a single view yields superior FID results; the reduced influx of information from neighboring views simplifies optimization for that view. However, this approach compromises mIoU, also reflecting less consistent generation, as depicted in Figure 10. Conversely, incorporating all views deteriorates performance across all metrics, potentially due to excessive information causing interference in cross-attention. Since each view has an intersection with both left and right views, attending to one view cannot guarantee consistency, especially for foreground objects, while attending to more views requires more computation. Thus, we opt for 2 attended views in our main paper, striking a balance between consistency and computational efficiency.
Appendix D More Experiments with 3D Object Detection
Table 6: Comparison about support for 3D object detection model (i.e., BEVFusion). MagicDrive generates 272×736 images for augmentation. Results are from tests on the nuScenes validation set.
Epoch Data CAM-Only CAM+LiDAR
mAP $\uparrow$ NDS $\uparrow$ mAP $\uparrow$ NDS $\uparrow$
0.5$\times$ w/o synthetic data 29.92 32.62 too few epochs too few epochs
w/ MagicDrive 32.88 (+2.96) 36.49 (+3.87)
1$\times$ w/o synthetic data 32.48 37.61 64.92 69.42
w/ MagicDrive 35.14 (+2.66) 39.63 (+2.02) 67.28 (+2.36) 70.14 (+0.72)
2$\times$ w/o synthetic data 35.16 40.49 67.87 71.12
w/ MagicDrive 35.40 (+0.24) 41.23 (+0.74) 68.15 (+0.28) 71.18 (+0.06)
In Table 6, we show additional experimental results on training 3D object detection models using synthetic data produced by MagicDrive. Given that BEVFusion utilizes a lightweight backbone (i.e., Swin-T (Liu et al., 2021)), model performance appears to plateau with training through 1-2$\times$ epochs (2$\times$: 20 for CAM-Only and 6 for CAM+LiDAR). Reducing epochs can mitigate this saturation, allowing more varied data to enhance the model’s perceptual capacity in both settings. This improvement is evident even when epochs for 3D object detection are further reduced to 0.5$\times$. Our MagicDrive accurately augments street-view images with the annotations. Future works may focus on annotation sampling and construction strategies for synthetic data augmentation.
Appendix E More Discussion
More future work.
Note MagicDrive-generated street views by now can only perform as augmented samples to train with real data. It is exciting to train perception models totally with generated data, which will be explored in the future. Meanwhile, more flexible usage of the generated street views beyond data augmentation, especially incorporation with generative pre-training (Chen et al., 2023a; Zhili et al., 2023), contrastive learning (Chen et al., 2021; Liu et al., 2022b) and DM-based perception models (Chen et al., 2022; Ji et al., 2023), is also an appealing future research direction.
Appendix F More Generation Results
We show some corner-case (Li et al., 2022) generations in Figure 11, and more generations in Figure 12-Figure 14.
Figure 11: Generation from MagicDrive with corner-case annotations.
Figure 12: Generation from MagicDrive with annotations from nuScenes validation set.
Figure 13: Generation from MagicDrive with annotations from nuScenes validation set.
Figure 14: Generation from MagicDrive with annotations from nuScenes validation set.